OpenStudy (anonymous):

Suppose that y varies inversely with x. Write a function that models the inverse function. x = 1 when y = 12

OpenStudy (radar):

Try y= 12/x

OpenStudy (radar):

Or maybe you should use f(y) = 12/x

OpenStudy (radar):

Then f(1) = 12/1 = 12

OpenStudy (radar):

f(2) = 12/2 = 6 Notice how x increases, y decreases, this is known as an inverse relation.

OpenStudy (anonymous):

thanks

OpenStudy (radar):

You're welcome.